# A Comprehensive Survey on Self-Supervised Learning for Recommendation

### v2

@@ 3.1.3. Contrastive Objective
  • InfoNCE-based Objective: InfoNCE (Oord et al., 2018), a variant of Noise Contrastive Estimation (Gutmann and Hyvärinen, 2010), has gained wide adoption as a learning objective in the field of contrastive learning for recommendation systems. The mathematical formulation of InfoNCE can be expressed as follows:
@@
  • JS-based Objective: In addition to using InfoNCE estimation for mutual information, the lower bound can also be estimated using the Jensen-Shannon (JS) divergence (Hjelm et al., 2018; Jing et al., 2023). The derived learning objective is akin to combining InfoNCE with a standard binary cross-entropy loss (Xia et al., 2021b), applied to positive pairs and negative pairs:
@@
    Here, $\sigma$ represents the sigmoid function used to normalize the output of the critic function. The main idea behind this optimization is to assign the label $1$ to positive pairs and 00 to negative pairs, thereby increasing the predicted value for positive pairs and enhancing the similarity between them.
  • Explicit Objective: Both InfoNCE-based and JS-based objectives maximize the estimated lower bound of mutual information to maximize mutual information itself, which is theoretically guaranteed. Additionally, there are explicit objectives, such as minimizing Mean Square Error or maximizing cosine similarity between samples within a positive pair, that directly align positive pairs. These objectives are referred to as explicit objectives.
3.2. Generative Learning in Recommendation
In generative self-supervised learning (GL), the primary objective is to maximize the likelihood estimation of the real data distribution. This allows the learned meaningful representations to capture the underlying structure and patterns in the data, which can then be utilized for downstream tasks. In our taxonomy, we consider two aspects to differentiate various recommendation methods with GL: generative learning paradigm and generation target.
Figure 3. Taxonomy of Generative Learning in Recommender Systems.
@@ 3.2.1. Learning Paradigm
In the context of recommendation, self-supervised methods employing generative learning can be classified into three paradigms: Masked Autoencoding, Variational Autoencoding and Denoised Diffusion.
@@
  • Masked Autoencoding: In masked autoencoders, the learning process follows a mask-reconstruction approach, where the model reconstructs complete data from partial observations (He et al., 2022). In recommender systems, data can be in item sequences or user-item interaction graphs, each requiring specific masking techniques. For item sequences, certain items are randomly masked and fed into a transformer model to encode representations and reconstruct masked item features (Sun et al., 2019). For graph data, masking occurs at the edge or node level. In GraphMAE (Hou et al., 2022a), node features are masked for reconstruction, while in MaskGAE (Li et al., 2022b), edges are masked. This masking practice is common in generative self-supervised learning recommender models.
@@
  • Variational Autoencoding: The variational autoencoder (Kingma and Welling, 2013a) is another generative approach that maximizes the likelihood estimation with theoretical guarantees. Typically, it involves mapping the input data to latent factors that follow a normal Gaussian distribution. Subsequently, the model reconstructs the input data based on sampled latent factors. Naturally, during the optimization process of VAE-based models, the learned latent factors are regularized using KL-divergence to enforce them to follow a Gaussian distribution. In self-supervised learning methods employing variational autoencoders, the latent factors may further be regulated through additional approaches, such as enhancing distinguishability through adversarial learning or contrastive learning (Wang et al., 2022e).
@@
  • Denoised Diffusion: Denoising Diffusion is a generative model that generates new data samples by reversing a noising process. In the forward process, Gaussian noise is added to the original data in multiple steps, creating a sequence of noisy versions. In the reverse process, the model learns to remove the noise from the noisy versions, recovering the original data step by step. The model is trained to denoise the data at each step, learning to capture minor changes in the complex generation process. The diffusion model has recently been introduced into the recommendation for processing data like graphs (Jiang et al., 2024) or user interactions (Wang et al., 2023d).
3.2.2. Generation Target
@@
What pattern of the data will be considered as the label for generation in generative learning is yet another thing that needs to be considered to bring meaningful auxiliary self-supervised signals. In general, the generation target can vary for different methods and also in different recommendation scenarios. For instance, in sequential recommendation, the generation target can be the item in the sequence (Sun et al., 2019; Sachdeva et al., 2019), in order to model the relationships within the item sequence. For recommendation with interaction graph, the generation target can be the node/edge in the graph (Xia et al., 2023; Li et al., 2023e), in order to capture high-level topological relevance in the graph.
@@ 3.3. Adversarial Learning in Recommendation
In the context of adversarial learning (AL) in recommender systems, the discriminator plays a crucial role in distinguishing between generated fake samples and real samples. Similar to generative learning, we propose our taxonomy that encompasses AL in recommender systems from both the learning paradigm and discrimination target perspectives:
@@ 3.3.1. Learning Paradigm
  • Non-differentiable AL Another approach involves discriminating the output of the recommender system, specifically the recommended items. However, since the recommended results are discrete, back-propagation becomes challenging, creating a non-differentiable situation where the gradient of the discriminator cannot be directly propagated to the generator (Wang et al., 2017; Zhao et al., 2021a). To address this, reinforcement learning with policy gradient is incorporated (Sutton et al., 1999). In this scenario, the generator acts as an agent that interacts with the environment by predicting the next item based on previously interacted items. The discriminator serves as the reward function, providing a reward signal to guide the generator’s learning. The rewards from the discriminator are defined to emphasize different factors influencing recommendation quality, and it is optimized to assign higher rewards to true samples over generated ones, guiding the generator to produce high-quality recommendations.
@@ 3.3.2. Discrimination Target
Different recommendation algorithms lead to the generation of different inputs by the generator, which are then fed to the discriminator for discrimination. This process aims to enhance the generator’s ability to generate high-quality content, thereby approaching the ground truth. The specific discrimination objectives are designed based on the particular recommendation task. For instance, in the sequential recommendation, the discriminator may be employed to discriminate generated next-items (Lv et al., 2021) or the entire sequence. In other domains, the discriminator may supervise the quality of generated data or optimize the hidden features of the model (Su et al., 2022; Zhang et al., 2021b).
@@ 4.1. General Collaborative Filtering
Table 1. A summary of SSL recommendation methods in General Collaborative Filtering (Part 1). For alphabets in ”Category”, CL means Contrastive Learning; GL means Generative Learning; AL means Adversarial Learning
@@
SGL (Wu et al., 2021b) 2021 SIGIR Data-based Natural InfoNCE-based
@@
DirectAU (Wang et al., 2022d) 2022 KDD Model-based Natural Explicit
SHT (Xia et al., 2022a) 2022 KDD Model-based Score-based Explicit
HCCF (Xia et al., 2022b) 2022 SIGIR Model-based Natural InfoNCE-based
RGCF (Tian et al., 2022b) 2022 SIGIR Data-based Natural InfoNCE-based
@@
AdvInfoNCE (Zhang et al., 2024b) 2023 NeurIPS Model-based Natural InfoNCE-based
@@
DCCF (Ren et al., 2023b) 2023 SIGIR Model-based Natural InfoNCE-based
@@
VGCL (Yang et al., 2023c) 2023 SIGIR Data-based Natural & Score-based InfoNCE-based
XSimGCL (Yu et al., 2023a) 2023 TKDE Feature-based Natural InfoNCE-based
@@
SelfCF (Zhou et al., 2023a) 2023 TORS Data & Feature-based Natural Explicit
@@
GL Method Venue Year Generative Learning Paradigm Generation Target
CVAE (Li and She, 2017) 2017 KDD Variational Autoencoding User Rating & Item Content
@@
MD-CVAE (Zhu and Chen, 2022) 2022 WWW Variational Autoencoding User Rating & Item Content
SE-VAE (Cho and Oh, 2022) 2022 WWW Variational Autoencoding User Rating to Items
CaD-VAE (Wang et al., 2023a) 2023 SIGIR Variational Autoencoding User Rating to Items
DiffRec (Wang et al., 2023d) 2023 SIGIR Denoising Diffusion User Rating to Items
GFormer (Li et al., 2023e) 2023 SIGIR Masked Autoencoding Edge in User-Item Graph
@@
ABinCF (Wang et al., 2019b) 2019 AAAI Differentiable Item Index
AugCF (Wang et al., 2019c) 2019 KDD Differentiable User-Item-Rating Tuple
@@
Conv-GCF (Yuan et al., 2020b) 2020 CIKM Differentiable Latent Interaction Map
@@ 4.1.2. Contrastive Learning in CF
  • View Creation in CL-based collaborative filtering covers diverse paradigms. (i) Data-based View Creation. Existing works employ different augmentation techniques to create views of the data for contrastive learning. For example, Liu et al.(Liu et al., 2021c) and methods like SGL(Wu et al., 2021b) and SelfCF (Zhou et al., 2023a) use edge/node dropout to augment the user-item graph. RGCF (Tian et al., 2022b) incorporates graph structure learning to generate two different graph views, such as a denoised graph and a diversity graph. Similar ideas are also utilized in EGLN (Yang et al., 2021b). AdaMCL (Zhu et al., 2023) and SGCCL (Li et al., 2023a) create user-user and item-item graphs based on the data and then encode views for users and items. LightGCL (Cai et al., 2023) created another user-item graph based on SVD decomposition that captures global relationship modeling. Recently, RecDCL (Zhang et al., 2024a) interpolates historical and current embeddings to obtain augmented views for batch contrastive learning (ii) Feature-based View Creation. In collaborative filtering, various feature-based augmentations involve adding noise to node features during model inference. Methods such as SimGCL (Yu et al., 2022) and XSimGCL (Yu et al., 2023a) introduce controllable random noise to node embeddings during message passing, thereby creating noised views of nodes. Another approach, RocSE (Ye et al., 2023a), employs embedding perturbation techniques inspired by adversarial attacks to generate different embedding views. (iii) Model-based View Creation. Different methods also employ various neural modules to create views. On one hand, SimpleX (Mao et al., 2021), DirectAU (Wang et al., 2022d), RecDCL-FCL (Zhang et al., 2024a), CGCL (He et al., 2023), AdvInfoNCE (Zhang et al., 2024c), uCTRL (Lee et al., 2023) and XSimGCL (Yu et al., 2023a) directly utilize the encoded user/item embeddings from models as views. On the other hand, some methods incorporate carefully designed neural modules to encode views from different perspectives. For example, EGLN (Yang et al., 2021b) and BiGI (Cao et al., 2021) focus on global representation fusion. GDCL (Zhang et al., 2022b) approximates diffusion models on graphs. HCCF (Xia et al., 2022b) and SHT (Xia et al., 2022a) employ hypergraph global encoding. DCCF (Ren et al., 2023b) explores intent disentanglement. NCL (Lin et al., 2022b) and CGCL (He et al., 2023) incorporate diverse neighbor view discovery. VGCL (Yang et al., 2023c) utilizes variational latent space sampling. DENS (Lai et al., 2023) incorporates item factor encoding. Recently, AdaGCL (Jiang et al., 2023) designed a graph generative model and a graph denoising model to create adaptive contrastive views.
@@
  • Pair Sampling in this scenario includes both natural sampling and also score-based sampling. (i) Natural Sampling. Most of the works in collaborative filtering follow natural sampling of positive samples for the anchor views to create positive pairs. There are three common situations for natural sampling in CF. Firstly, when the model encodes multiple views for one user/item, any two of these views can form a pair (Wu et al., 2021b; Cai et al., 2023; Yu et al., 2022; Xia et al., 2022b; Ren et al., 2023b). Secondly, the views from the interacted user and item can form a pair (Wang et al., 2022d; Lee et al., 2023; Zhang et al., 2024c). Lastly, by considering the relevant/irrelevant factor relationship with user-item interaction, pairs of views can be naturally formed (Lai et al., 2023). (ii) Score-based Sampling. SHT calculates edge solidity scores and samples pairs of edges for learning. VGCL clusters users/items based on embedding similarity and pairs users from the same cluster.
@@
  • Contrastive Objective utilizes InfoNCE-based or JS-based objectives to optimize mutual information. Notably, AdvInfoNCE improves the weighting of negative pairs in InfoNCE by incorporating hardness scores learned through a min-max game. Some methods explicitly design loss functions to enhance the similarity between positive pairs. This can be achieved through cosine similarity optimization (Mao et al., 2021; Zhou et al., 2023a; Lai et al., 2023), alignment/uniformity regularization (Wang et al., 2022d; Lee et al., 2023; Zhang et al., 2024a), or by explicitly regulating the score difference between pairs (Xia et al., 2022a).
4.1.3. Generative Learning in CF

### v3

@@ 3.1.3. Contrastive Objective
  • InfoNCE-based Objective. InfoNCE (Oord et al., 2018), a variant of Noise Contrastive Estimation (Gutmann et al., 2010), has gained wide adoption as a learning objective in the field of contrastive learning for recommendation systems. The mathematical formulation of InfoNCE can be expressed as follows:
@@
  • JS-based Objective. In addition to using InfoNCE estimation for mutual information, the lower bound can also be estimated using the Jensen-Shannon (JS) divergence (Hjelm et al., 2018; Jing et al., 2023). The derived learning objective is akin to combining InfoNCE with a standard binary cross-entropy loss (Xia et al., 2021b), applied to positive pairs and negative pairs:
@@
    Here, $\sigma$ represents the sigmoid function used to normalize the output of the critic function. The main idea behind this optimization is to assign the label $1$ to positive pairs and 0 to negative pairs, thereby increasing the predicted value for positive pairs and enhancing the similarity between them.
  • Explicit Objective. Both InfoNCE-based and JS-based objectives maximize the estimated lower bound of mutual information to maximize mutual information itself, which is theoretically guaranteed. Additionally, there are explicit objectives, such as minimizing Mean Square Error or maximizing cosine similarity between samples within a positive pair, that directly align positive pairs. These objectives are referred to as explicit objectives.
3.2. Reconstructive Learning in Recommendation
In reconstructive self-supervised learning (RL), the primary objective is to maximize the likelihood estimation of the real data distribution. This allows the learned meaningful representations to capture the underlying patterns in the data, which can then be utilized for downstream tasks. In our taxonomy (as illustrated in Figure 3), we consider two aspects to differentiate various recommendation methods with RL: reconstructive learning paradigm and generation target.
Figure 3. Taxonomy of Reconstructive Learning in Recommender Systems.
@@ 3.2.1. Learning Paradigm
In the context of recommendation, recent SSL methods employing reconstructive learning can be classified into three paradigms: Masked Autoencoding, Variational Autoencoding and Denoised Diffusion.
@@
  • Masked Autoencoding. In masked autoencoders, the learning process follows a mask-reconstruction approach, where the model reconstructs complete data from partial observations (He et al., 2022). In recommender systems, data can be in item sequences or user-item interaction graphs, each requiring specific masking techniques. For item sequences, certain items are randomly masked and fed into a transformer model to encode representations and reconstruct masked item features (Sun et al., 2019). For graph data, masking occurs at the edge or node level. In GraphMAE (Hou et al., 2022), node features are masked for reconstruction, while in MaskGAE (Li et al., 2023c), edges are masked. This masking practice is common in reconstructive self-supervised learning recommender models.
@@
  • Variational Autoencoding. The variational autoencoder (Kingma and Welling, 2013) is another reconstructive approach that maximizes the likelihood estimation with theoretical guarantees. Typically, it involves mapping the input data to latent factors that follow a normal Gaussian distribution. Subsequently, the model reconstructs the input data based on sampled latent factors. Naturally, during the optimization process of VAE-based models, the learned latent factors are regularized using KL-divergence to enforce them to follow a Gaussian distribution. In self-supervised learning methods employing variational autoencoders, the latent factors may further be regulated through additional approaches, such as enhancing distinguishability through adversarial learning or contrastive learning (Wang et al., 2022d).
@@
  • Denoised Diffusion. Denoising diffusion is a generative model that reconstructs the data samples by reversing a noising process. In the forward process, Gaussian noise is added to the original data in multiple steps, creating a sequence of noisy versions. In the reverse process, the model learns to remove the noise from the noisy versions, recovering the original data step by step. The model is trained to denoise the data at each step, learning to capture minor changes in the complex generation process. The diffusion model has recently been introduced into the recommendation for processing data like graphs (Jiang et al., 2024a) or user interactions (Wang et al., 2023e).
3.2.2. Reconstruction Target
@@
What pattern of the data will be considered as the label for reconstruction in reconstructive learning is yet another thing that needs to be considered to bring meaningful auxiliary self-supervised signals. In general, the target can vary for different methods and also in different recommendation scenarios. For instance, in sequential recommendation, the reconstruction target can be the item in the sequence (Sun et al., 2019; Sachdeva et al., 2019), in order to model the relationships within the item sequence. For recommendation with interaction graph, the target can be the node/edge in the graph (Xia et al., 2023; Li et al., 2023b), in order to capture high-level topological relevance in the graph.
@@ 3.3. Adversarial Learning in Recommendation
In the context of adversarial learning (AL) for recommender systems, the discriminator serves as a critical component that distinguishes between generated fake samples and authentic real samples. Drawing parallels to reconstructive learning, we establish our taxonomy (as depicted in Figure 4) that characterizes AL-based recommendation methods from two key perspectives: the learning paradigm and the discrimination target.
@@ 3.3.1. Learning Paradigm
  • Non-differentiable AL Another approach involves discriminating the output of the recommender system, specifically the recommended items. However, the discrete nature of these results complicates back-propagation, making it non-differentiable and preventing the gradient from being directly propagated to the generator (Wang et al., 2017; Zhao et al., 2021b). To address this, reinforcement learning with policy gradients is used (Sutton et al., 1999). Here, the generator acts as an agent that predicts the next item based on prior interactions, while the discriminator serves as the reward function, guiding the generator’s learning. The rewards emphasize various factors affecting recommendation quality, optimizing to assign higher rewards to true samples compared to generated ones, thus encouraging high-quality recommendations.
@@ 3.3.2. Discrimination Target
Different recommendation algorithms produce varied inputs for the generator, which are then evaluated by the discriminator. This process enhances the generator’s ability to create high-quality content that aligns with the ground truth. Discrimination objectives are tailored to specific recommendation tasks. For example, in sequential recommendation, the discriminator may assess generated next-items (Lv et al., 2021) or entire sequences. In other areas, it may oversee data quality or optimize the model’s hidden features (Su et al., 2022; Zhang et al., 2021b).
3.4. Analysis of Self-supervised Learning Techniques
In the realm of self-supervised learning (SSL) for recommender systems, various techniques-namely adversarial, reconstructive, and contrastive learning-offer distinct advantages tailored to specific challenges. Here, we provide an in-depth analysis of the merits of each method and the trends in the development of SSL techniques for recommendation:
  • Merits of Diverse SSL Techniques. Adversarial learning enhances model robustness by using a discriminator to distinguish between real and generated samples, which is crucial for maintaining data quality. Consequently, this method encourages the generator to produce high-fidelity recommendations, while the integration of reinforcement learning provides adaptive feedback for further refinement. In contrast, reconstructive learning focuses on capturing the underlying data distribution through techniques such as variational autoencoders and masked autoencoders. This approach is effective for modeling complex structures and enables data reconstruction, thereby facilitating augmentation that improves overall model performance. Meanwhile, contrastive learning prioritizes maximizing agreement between different views of the same data, effectively distinguishing positive and negative pairs. As a result, this technique captures nuanced relationships and enhances generalization, particularly in scenarios where labeled data is sparse. In summary, while adversarial and reconstructive techniques improve data quality and representation, contrastive learning excels at extracting meaningful insights from unlabeled datasets. Hence, the choice of techniques usually align with the specific requirements of the task.
  • Evolution in SSL-based Recommendation. The evolution of self-supervised learning (SSL) techniques in recommendation has generally paralleled advancements in self-supervised methods within deep learning. Prior to 2021, SSL algorithms primarily relied on variational autoencoders (Li et al., 2017; Liang et al., 2018) and adversarial learning (Manotumruksa et al., 2020; Lu et al., 2021), significantly enhancing model performance and robustness. Following 2021, two major trends emerged: first, the development of masked autoencoding (Hou et al., 2022) and reconstructive learning (Sun et al., 2019) led to effective applications in sequential recommendation; second, the rise and theoretical refinement of contrastive learning (Wu et al., 2021a; Xia et al., 2022a) resulted in numerous research studies focused on contrastive-based recommendation methods. Recently, the introduction of diffusion models (Wang et al., 2023e) has emerged as a new trend in the study of SSL-based recommendation.
@@ 4.1. General Collaborative Filtering
Table 1. A summary of SSL recommendation methods in General Collaborative Filtering (Part 1). For alphabets in ”Category”, CL means Contrastive Learning; RL means Reconstructive Learning; AL means Adversarial Learning
@@
SGL (Wu et al., 2021a) 2021 SIGIR Data-based Natural InfoNCE-based
@@
DirectAU (Wang et al., 2022a) 2022 KDD Model-based Natural Explicit
SHT (Xia et al., 2022b) 2022 KDD Model-based Score-based Explicit
HCCF (Xia et al., 2022a) 2022 SIGIR Model-based Natural InfoNCE-based
RGCF (Tian et al., 2022a) 2022 SIGIR Data-based Natural InfoNCE-based
@@
CCL (Zhou et al., 2023a) 2023 CIKM Data-based Score-based InfoNCE-based
@@
AdvInfoNCE (Zhang et al., 2023a) 2023 NeurIPS Model-based Natural InfoNCE-based
@@
DCCF (Ren et al., 2023a) 2023 SIGIR Model-based Natural InfoNCE-based
@@
VGCL (Yang et al., 2023b) 2023 SIGIR Data-based Natural & Score-based InfoNCE-based
XSimGCL (Yu et al., 2023b) 2023 TKDE Feature-based Natural InfoNCE-based
@@
SelfCF (Zhou et al., 2023c) 2023 TORS Data & Feature-based Natural Explicit
@@
DivGCL (Gong et al., 2025) 2025 AAAI Model-based Natural InfoNCE-based
@@
RL Method Venue Year Reconstructive Learning Paradigm Reconstruction Target
CVAE (Li et al., 2017) 2017 KDD Variational Autoencoding User Rating & Item Content
@@
MD-CVAE (Zhu et al., 2022) 2022 WWW Variational Autoencoding User Rating & Item Content
SE-VAE (Cho et al., 2022) 2022 WWW Variational Autoencoding User Rating to Items
CaD-VAE (Wang et al., 2023c) 2023 SIGIR Variational Autoencoding User Rating to Items
DiffRec (Wang et al., 2023e) 2023 SIGIR Denoising Diffusion User Rating to Items
GFormer (Li et al., 2023b) 2023 SIGIR Masked Autoencoding Edge in User-Item Graph
@@
ABinCF (Wang et al., 2019a) 2019 AAAI Differentiable Item Index
AugCF (Wang et al., 2019b) 2019 KDD Differentiable User-Item-Rating Tuple
ACAE (Yuan et al., 2019) 2019 SIGIR Differentiable User Purchase Vector
@@
Conv-GCF (Yuan et al., 2020a) 2020 CIKM Differentiable Latent Interaction Map
@@ 4.1.2. Contrastive Learning in CF
  • View Creation in CL-based CF covers diverse paradigms. (i) Data-based View Creation. Existing works employ different augmentation techniques to create views of the data for contrastive learning. For example, Liu et al.(Liu et al., 2021c) and methods like SGL(Wu et al., 2021a) and SelfCF (Zhou et al., 2023c) use edge/node dropout to augment the user-item graph. RGCF (Tian et al., 2022a) incorporates graph structure learning to generate two different graph views, such as a denoised graph and a diversity graph. Similar ideas are also utilized in EGLN (Yang et al., 2021b). AdaMCL (Zhu et al., 2023) and SGCCL (Li et al., 2023a) create user-user and item-item graphs based on the data and then encode views for users and items. LightGCL (Cai et al., 2023) created another user-item graph based on SVD decomposition that captures global relationship modeling. Recently, RecDCL (Zhang et al., 2024a) interpolates historical and current embeddings to obtain augmented views for batch contrastive learning (ii) Feature-based View Creation. In collaborative filtering, various feature-based augmentations involve adding noise to node features during model inference. Methods such as SimGCL (Yu et al., 2022) and XSimGCL (Yu et al., 2023b) introduce controllable random noise to node embeddings during message passing, thereby creating noised views of nodes. Another approach, RocSE (Ye et al., 2023a), employs embedding perturbation techniques inspired by adversarial attacks to generate different embedding views. (iii) Model-based View Creation. Different methods also employ various neural modules to create views. On one hand, SimpleX (Mao et al., 2021), DirectAU (Wang et al., 2022a), RecDCL-FCL (Zhang et al., 2024a), CGCL (He et al., 2023), AdvInfoNCE (Zhang et al., 2023a), uCTRL (Lee et al., 2023) and XSimGCL (Yu et al., 2023b) directly utilize the encoded user/item embeddings from models as views. On the other hand, some methods incorporate carefully designed neural modules to encode views from different perspectives. For example, EGLN (Yang et al., 2021b) and BiGI (Cao et al., 2021) focus on global representation fusion. GDCL (Zhang et al., 2022b) approximates diffusion models on graphs. HCCF (Xia et al., 2022a) and SHT (Xia et al., 2022b) employ hypergraph global encoding. DCCF (Ren et al., 2023a) explores intent disentanglement. NCL (Lin et al., 2022b) and CGCL (He et al., 2023) incorporate diverse neighbor view discovery. VGCL (Yang et al., 2023b) utilizes variational latent space sampling. DENS (Lai et al., 2023) incorporates item factor encoding. AdaGCL (Jiang et al., 2023) designed a graph model and a denoising model to create adaptive contrastive views.
@@
  • Pair Sampling in this scenario includes both natural sampling and also score-based sampling. (i) Natural Sampling. Most of the works in collaborative filtering follow natural sampling of positive samples for the anchor views to create positive pairs. There are three common situations for natural sampling in CF. Firstly, when the model encodes multiple views for one user/item, any two of these views can form a pair (Wu et al., 2021a; Cai et al., 2023; Yu et al., 2022; Xia et al., 2022a; Ren et al., 2023a). Secondly, the views from the interacted user and item can form a pair (Wang et al., 2022a; Lee et al., 2023; Zhang et al., 2023a). Lastly, by considering the relevant/irrelevant factor relationship with user-item interaction, pairs of views can be naturally formed (Lai et al., 2023). (ii) Score-based Sampling. SHT calculates edge solidity scores and samples pairs of edges for learning. VGCL clusters users/items based on embedding similarity and pairs users from the same cluster.
@@
  • Contrastive Objective utilizes InfoNCE-based or JS-based objectives to optimize mutual information. Notably, AdvInfoNCE improves the weighting of negative pairs in InfoNCE by incorporating hardness scores learned through a min-max game. Some methods explicitly design loss functions to enhance the similarity between positive pairs. This can be achieved through cosine similarity optimization (Mao et al., 2021; Zhou et al., 2023c; Lai et al., 2023), alignment/uniformity regularization (Wang et al., 2022a; Lee et al., 2023; Zhang et al., 2024a), or by explicitly regulating the score difference between pairs (Xia et al., 2022b).
4.1.3. Reconstructive Learning in CF